# First Comparative Exoplanetology Within a Transiting Multi-planet System: Comparing the atmospheres of V1298 Tau b and c

### v1

@@ 6 Acknowledgements
J.M.D acknowledges support from the Amsterdam Academic Alliance (AAA) Program, and the European Research Council (ERC) European Union’s Horizon 2020 research and innovation program (grant agreement no. 679633; Exo-Atmos). This work is part of the research program VIDI New Frontiers in Exoplanetary Climatology with project number 614.001.601, which is (partly) financed by the Dutch Research Council (NWO). Y.K. acknowledges the support from JSPS KAKENHI Grant Numbers JP21K13984 and 22H05150.
7 Softwares used
numpy, scipy, matplotlib, emcee, lmfit, pandas, ATMO, PetitRadtrans
@@ Appendix A Description of raw data reduction
We start our analysis from the ‘ima.fits’ files downloaded from MAST archive. During each exposure, the detector is read multiple times, without flushing out the accumulated charge (non-destructive reads). We create sub-exposures by subtracting consecutive non-destructive reads for each exposure which are reduced separately. This technique has been widely applied in past observations with HST (e.g, see Kreidberg et al. 2014; Arcangeli et al. 2018; Jacobs et al. 2022). We apply a wavelength dependent flat-field correction and flag bad pixels with data quality DQ=4, 32, or 512 by calwf3. We apply a local median filter to identify cosmic rays and clip pixels which are more than five median deviations from the median. To account for the dispersion direction drift of the spectrum (Wakeford et al. 2016) we use the first exposure of a visit as a template and shift the spectrum for an exposure along the dispersion direction to match the template. The maximum shift that we require is 0.1 pixels. We apply optimal extraction algorithm (Horne 1986) on each subexposure to maximize signal-to-noise ratio.
@@ Appendix B Light curve fitting
HST WFC3 light curves often exhibit strong systematics like charge-trapping ramps and visit-long slopes (Kreidberg et al. 2014; Deming et al. 2013; Tsiaras et al. 2016; Wakeford et al. 2016). Common practice involves discarding the first orbit due to significantly larger systematics (See Figure 4) compared to the rest of the visit (e.g see, Wakeford et al. 2013; Arcangeli et al. 2018; Jacobs et al. 2022; Barat et al. 2024). However, in this case, we could not remove from the analysis because the transit occurred two hours earlier than expected from the linear ephemeris (Feinstein et al. 2022) due to large TTVs and we could only secure one orbit prior to ingress.
@@
The first-orbit systematics are strongly wavelength dependent (Zhou et al. 2020, 2017) and cannot be corrected for using common-mode approaches, which had been previously applied to similar observations of V1298 tau b (Barat et al. 2024). Therefore, to handle these systematics, we used the physically motivated RECTE charge-trapping model (Zhou et al. 2017), which fits the first orbit ramp better than other methods (Berta et al. 2012; Kreidberg et al. 2014). We fixed the mid-transit time using simultaneous TESS observations (Feinstein et al. 2022). V1298 Tau shows strong rotational variability, with an average amplitude of 2 and a rotational period of 2.83 days (David et al. 2019b). We modeled it with a second-order polynomial for the out-of-transit baseline. See Appendix C for more discussion on the baseline modelling. The planetary transit signal was modelled using batman (Kreidberg 2015).
@@
For the white light curve (1.12-1.6 m), we fit the RECTE parameters and the baseline polynomial coefficients, along with the planet’s transit depth and semi-major axis. Other orbital parameters were fixed to literature values (Feinstein et al. 2022; Sikora et al. 2023; David et al. 2019a). The RECTE parameters constitute 4 parameters: the number of initially filled fast/slow charge traps and the number of fast/slow charge traps that are filled in between orbits. Due to the lack of ingress and egress data, our limb darkening precision was impaired. However, by using the limb darkening parameter retrieved from the observations of planet b (Barat et al. 2024), we were able to significantly reduce the uncertainties on the transit spectrum (Figure 8). This highlights the advantage and uniqueness of studying multiple planets in the same system. For a discussion on the limb darkening of V1298 Tau see Appendix 4.2. We also independently measure the stellar density from the HST white light curves which is discussed in Appendix 4.1.
@@
The fit was performed using emcee (Foreman-Mackey et al. 2013). The posterior distribution from these fits is shown in Supplementary Figure 5 and the best-fit transit model and residuals, are shown in Supplementary Figure 4 (upper panel).
@@
Spectroscopic light curves were analyzed similarly by dividing the stellar spectra into 7-pixel bins, resulting in 17 channels (identical to V1298 Tau b). We fixed the semi-major axis from the white light curve fits and fit the RECTE parameters and baseline for each bin, and the transit depth. We fix the wavelength dependent limb-darkening coefficients to values derived in Barat et al. (2024). The detrended spectroscopic light curves and residuals are shown in Supplementary Figure 6.
Appendix C Modelling the baseline

### v2

@@ Acknowledgements.
J.M.D acknowledges support from the Amsterdam Academic Alliance (AAA) Program, and the European Research Council (ERC) European Union’s Horizon 2020 research and innovation program (grant agreement no. 679633; Exo-Atmos). This work is part of the research program VIDI New Frontiers in Exoplanetary Climatology with project number 614.001.601, which is (partly) financed by the Dutch Research Council (NWO). Y.K. acknowledges the support from JSPS KAKENHI Grant Numbers 21K13984, 22H05150, and 23H01224. We acknowledge publicly available open source softwares which have been used for this work: numpy, scipy, matplotlib, emcee, lmfit, pandas, ATMO, PetitRadtransReferences
- Ahrer et al. (2023) Ahrer, E.-M., Stevenson, K. B., Mansfield, M., et al. 2023, Nature, 614, 653
- Alderson et al. (2023) Alderson, L., Wakeford, H. R., Alam, M. K., et al. 2023, Nature, 614, 664
- Arcangeli et al. (2018) Arcangeli, J., Désert, J.-M., Line, M. R., et al. 2018, The Astrophysical Journal, 855, L30
- Arcangeli et al. (2019) Arcangeli, J., Désert, J.-M., Parmentier, V., et al. 2019, A&A, 625, A136
- Arcangeli et al. (2021) Arcangeli, J., Désert, J. M., Parmentier, V., Tsai, S. M., & Stevenson, K. B. 2021, A&A, 646, A94
- August et al. (2023) August, P. C., Bean, J. L., Zhang, M., et al. 2023, The Astrophysical Journal Letters, 953, L24
- Baeyens et al. (2022) Baeyens, R., Konings, T., Venot, O., Carone, L., & Decin, L. 2022, MNRAS, 512, 4877
- Barat et al. (2024) Barat, S., Désert, J.-M., Vazan, A., et al. 2024, Nature Astronomy [arXiv:2312.16924]
- Barclay et al. (2021) Barclay, T., Kostov, V. B., Colón, K. D., et al. 2021, AJ, 162, 300
- Baxter et al. (2020) Baxter, C., Désert, J.-M., Parmentier, V., et al. 2020, A&A, 639, A36
- Baxter et al. (2021) Baxter, C., Désert, J.-M., Tsai, S.-M., et al. 2021, A&A, 648, A127
- Bean et al. (2010) Bean, J. L., Miller-Ricci Kempton, E., & Homeier, D. 2010, Nature, 468, 669
- Bean et al. (2021) Bean, J. L., Raymond, S. N., & Owen, J. E. 2021, Journal of Geophysical Research (Planets), 126, e06639
- Bell et al. (2023) Bell, T. J., Welbanks, L., Schlawin, E., et al. 2023, Nature, 623, 709
- Benneke et al. (2024) Benneke, B., Roy, P.-A., Coulombe, L.-P., et al. 2024, arXiv e-prints, arXiv:2403.03325
- Benz et al. (2014) Benz, W., Ida, S., Alibert, Y., Lin, D., & Mordasini, C. 2014, in Protostars and Planets VI, ed. H. Beuther, R. S. Klessen, C. P. Dullemond, & T. Henning, 691–713
- Berta et al. (2012) Berta, Z. K., Charbonneau, D., Désert, J.-M., et al. 2012, ApJ, 747, 35
- Bitsch et al. (2019) Bitsch, B., Raymond, S. N., & Izidoro, A. 2019, A&A, 624, A109
- Blunt et al. (2023) Blunt, S., Carvalho, A., David, T. J., et al. 2023, AJ, 166, 62
- Bodenheimer & Pollack (1986) Bodenheimer, P. & Pollack, J. B. 1986, Icarus, 67, 391
- Bodenheimer et al. (2018) Bodenheimer, P., Stevenson, D. J., Lissauer, J. J., & D’Angelo, G. 2018, ApJ, 868, 138
- Boss (1997) Boss, A. P. 1997, Science, 276, 1836
- Bourque et al. (2021) Bourque, M., Espinoza, N., Filippazzo, J., et al. 2021, The Exoplanet Characterization Toolkit (ExoCTK)
- Brande et al. (2024) Brande, J., Crossfield, I. J. M., Kreidberg, L., et al. 2024, The Astrophysical Journal Letters, 961, L23
- Brems et al. (2019) Brems, S. S., Kürster, M., Trifonov, T., Reffert, S., & Quirrenbach, A. 2019, A&A, 632, A37
- Carter et al. (2020) Carter, A. L., Nikolov, N., Sing, D. K., et al. 2020, MNRAS, 494, 5449
- Coulombe et al. (2023) Coulombe, L.-P., Benneke, B., Challener, R., et al. 2023, Nature, 620, 292
- Damasso et al. (2023) Damasso, M., Scandariato, G., Nascimbeni, V., et al. 2023, A&A, 680, A8
- David et al. (2019a) David, T. J., Cody, A. M., Hedges, C. L., et al. 2019a, AJ, 158, 79
- David et al. (2016) David, T. J., Hillenbrand, L. A., Petigura, E. A., et al. 2016, Nature, 534, 658
- David et al. (2019b) David, T. J., Petigura, E. A., Luger, R., et al. 2019b, ApJ, 885, L12
- Deming et al. (2013) Deming, D., Wilkins, A., McCullough, P., et al. 2013, ApJ, 774, 95
- Désert et al. (2011a) Désert, J.-M., Bean, J., Miller-Ricci Kempton, E., et al. 2011a, ApJ, 731, L40
- Désert et al. (2011b) Désert, J.-M., Charbonneau, D., Demory, B.-O., et al. 2011b, ApJS, 197, 14
- Dodson-Robinson et al. (2009) Dodson-Robinson, S. E., Veras, D., Ford, E. B., & Beichman, C. A. 2009, The Astrophysical Journal, 707, 79
- Drummond et al. (2016) Drummond, B., Tremblin, P., Baraffe, I., et al. 2016, A&A, 594, A69
- Duvvuri et al. (2023) Duvvuri, G. M., Cauley, P. W., Aguirre, F. C., et al. 2023, The Astronomical Journal, 166, 196
- Emsenhuber et al. (2023) Emsenhuber, A., Mordasini, C., & Burn, R. 2023, European Physical Journal Plus, 138, 181
- Erkaev et al. (2007) Erkaev, N. V., Kulikov, Y. N., Lammer, H., et al. 2007, A&A, 472, 329
- Evans et al. (2018) Evans, T. M., Sing, D. K., Goyal, J. M., et al. 2018, AJ, 156, 283
- Feinstein et al. (2022) Feinstein, A. D., David, T. J., Montet, B. T., et al. 2022, The Astrophysical Journal Letters, 925, L2
- Feinstein et al. (2021) Feinstein, A. D., Montet, B. T., Johnson, M. C., et al. 2021, AJ, 162, 213
- Feinstein et al. (2023) Feinstein, A. D., Radica, M., Welbanks, L., et al. 2023, Nature, 614, 670
- Finociety et al. (2023) Finociety, B., Donati, J. F., Cristofari, P. I., et al. 2023, MNRAS, 526, 4627
- Foreman-Mackey et al. (2013) Foreman-Mackey, D., Hogg, D. W., Lang, D., & Goodman, J. 2013, PASP, 125, 306
- Fortney et al. (2013) Fortney, J. J., Mordasini, C., Nettelmann, N., et al. 2013, 775, 80
- Fortney et al. (2020) Fortney, J. J., Visscher, C., Marley, M. S., et al. 2020, AJ, 160, 288
- Fu et al. (2017) Fu, G., Deming, D., Knutson, H., et al. 2017, The Astrophysical Journal Letters, 847, L22
- Fulton & Petigura (2018) Fulton, B. J. & Petigura, E. A. 2018, AJ, 156, 264
- Fulton et al. (2017) Fulton, B. J., Petigura, E. A., Howard, A. W., et al. 2017, The Astronomical Journal, 154, 109
- Gagnebin et al. (2024) Gagnebin, A., Mukherjee, S., Fortney, J. J., & Batalha, N. E. 2024, ApJ, 969, 86
- Gao et al. (2020) Gao, P., Thorngren, D. P., Lee, E. K. H., et al. 2020, Nature Astronomy, 4, 951
- Gao & Zhang (2020) Gao, P. & Zhang, X. 2020, ApJ, 890, 93
- Garhart et al. (2020) Garhart, E., Deming, D., Mandell, A., et al. 2020, AJ, 159, 137
- Ginzburg et al. (2018) Ginzburg, S., Schlichting, H. E., & Sari, R. 2018, MNRAS, 476, 759
- Goyal et al. (2018) Goyal, J. M., Mayne, N., Sing, D. K., et al. 2018, MNRAS, 474, 5158
- Goyal et al. (2019) Goyal, J. M., Wakeford, H. R., Mayne, N. J., et al. 2019, MNRAS, 482, 4503
- Grant et al. (2023) Grant, D., Lewis, N. K., Wakeford, H. R., et al. 2023, The Astrophysical Journal Letters, 956, L32
- GRAVITY Collaboration et al. (2019) GRAVITY Collaboration, Lacour, S., Nowak, M., et al. 2019, A&A, 623, L11
- Gupta & Schlichting (2019) Gupta, A. & Schlichting, H. E. 2019, Monthly Notices of the Royal Astronomical Society, 487, 24
- Heng & Kitzmann (2017) Heng, K. & Kitzmann, D. 2017, MNRAS, 470, 2972
- Horne (1986) Horne, K. 1986, PASP, 98, 609
- Jacobs et al. (2022) Jacobs, B., Désert, J. M., Pino, L., et al. 2022, A&A, 668, L1
- Johnson et al. (2022) Johnson, M. C., David, T. J., Petigura, E. A., et al. 2022, AJ, 163, 247
- Kajtazi et al. (2023) Kajtazi, K., Petit, A. C., & Johansen, A. 2023, A&A, 669, A44
- Kawashima & Ikoma (2018) Kawashima, Y. & Ikoma, M. 2018, The Astrophysical Journal, 853, 7
- Kawashima & Ikoma (2019) Kawashima, Y. & Ikoma, M. 2019, ApJ, 877, 109
- Keating et al. (2019) Keating, D., Cowan, N. B., & Dang, L. 2019, Nature Astronomy, 3, 1092
- Kempton et al. (2023) Kempton, E. M. R., Zhang, M., Bean, J. L., et al. 2023, Nature, 620, 67
- Khare et al. (1984) Khare, B. N., Sagan, C., Arakawa, E. T., et al. 1984, Icarus, 60, 127
- Knutson et al. (2014) Knutson, H. A., Benneke, B., Deming, D., & Homeier, D. 2014, Nature, 505, 66
- Koen (2016) Koen, C. 2016, MNRAS, 463, 4383
- Kostogryz et al. (2024) Kostogryz, N. M., Shapiro, A. I., Witzke, V., et al. 2024, arXiv e-prints, arXiv:2403.00118
- Kreidberg (2015) Kreidberg, L. 2015, PASP, 127, 1161
- Kreidberg et al. (2014) Kreidberg, L., Bean, J. L., Désert, J.-M., et al. 2014, Nature, 505, 69
- Kreidberg et al. (2022) Kreidberg, L., Mollière, P., Crossfield, I. J. M., et al. 2022, AJ, 164, 124
- Kubyshkina et al. (2020) Kubyshkina, D., Vidotto, A. A., Fossati, L., & Farrell, E. 2020, MNRAS, 499, 77
- Lambrechts et al. (2019) Lambrechts, M., Morbidelli, A., Jacobson, S. A., et al. 2019, A&A, 627, A83
- Landman et al. (2024) Landman, R., Stolker, T., Snellen, I. A. G., et al. 2024, A&A, 682, A48
- Lee & Chiang (2015) Lee, E. J. & Chiang, E. 2015, The Astrophysical Journal, 811, 41
- Lee & Chiang (2016) Lee, E. J. & Chiang, E. 2016, ApJ, 817, 90
- Lee et al. (2014) Lee, E. J., Chiang, E., & Ormel, C. W. 2014, ApJ, 797, 95
- Lewis et al. (2020) Lewis, N. K., Wakeford, H. R., MacDonald, R. J., et al. 2020, ApJ, 902, L19
- Libby-Roberts et al. (2020a) Libby-Roberts, J. E., Berta-Thompson, Z. K., Désert, J.-M., et al. 2020a, AJ, 159, 57
- Libby-Roberts et al. (2020b) Libby-Roberts, J. E., Berta-Thompson, Z. K., Désert, J.-M., et al. 2020b, AJ, 159, 57
- Lim et al. (2023) Lim, O., Benneke, B., Doyon, R., et al. 2023, ApJ, 955, L22
- Linder et al. (2019) Linder, E. F., Mordasini, C., Mollière, P., et al. 2019, A&A, 623, A85
- Luhman (2018) Luhman, K. L. 2018, AJ, 156, 271
- Madhusudhan et al. (2023) Madhusudhan, N., Sarkar, S., Constantinou, S., et al. 2023, ApJ, 956, L13
- Mann et al. (2016) Mann, A. W., Newton, E. R., Rizzuto, A. C., et al. 2016, The Astronomical Journal, 152, 61
- Mann et al. (2022) Mann, A. W., Wood, M. L., Schmidt, S. P., et al. 2022, AJ, 163, 156
- Marois et al. (2010) Marois, C., Zuckerman, B., Konopacky, Q. M., Macintosh, B., & Barman, T. 2010, Nature, 468, 1080
- May et al. (2023) May, E. M., MacDonald, R. J., Bennett, K. A., et al. 2023, ApJ, 959, L9
- Miller-Ricci Kempton et al. (2012) Miller-Ricci Kempton, E., Zahnle, K., & Fortney, J. J. 2012, ApJ, 745, 3
- Mollière et al. (2019) Mollière, P., Wardenier, J. P., van Boekel, R., et al. 2019, A&A, 627, A67
- Moran et al. (2023) Moran, S. E., Stevenson, K. B., Sing, D. K., et al. 2023, ApJ, 948, L11
- Mordasini (2018) Mordasini, C. 2018, in Handbook of Exoplanets, ed. H. J. Deeg & J. A. Belmonte, 143
- Morley et al. (2013) Morley, C. V., Fortney, J. J., Kempton, E. M. R., et al. 2013, ApJ, 775, 33
- Nasedkin et al. (2024) Nasedkin, E., Mollière, P., Lacour, S., et al. 2024, A&A, 687, A298
- Öberg et al. (2011) Öberg, K. I., Murray-Clay, R., & Bergin, E. A. 2011, ApJ, 743, L16
- Oh et al. (2017) Oh, S., Price-Whelan, A. M., Hogg, D. W., Morton, T. D., & Spergel, D. N. 2017, VizieR Online Data Catalog, J/AJ/153/257
- Ormel et al. (2021) Ormel, C. W., Vazan, A., & Brouwers, M. G. 2021, A&A, 647, A175
- Owen & Wu (2013) Owen, J. E. & Wu, Y. 2013, ApJ, 775, 105
- Owen & Wu (2017) Owen, J. E. & Wu, Y. 2017, The Astrophysical Journal, 847, 29
- Pollack et al. (1996) Pollack, J. B., Hubickyj, O., Bodenheimer, P., et al. 1996, Icarus, 124, 62
- Pont et al. (2008) Pont, F., Knutson, H., Gilliland, R. L., Moutou, C., & Charbonneau, D. 2008, MNRAS, 385, 109
- Rackham et al. (2019) Rackham, B. V., Apai, D., & Giampapa, M. S. 2019, AJ, 157, 96
- Radica et al. (2023) Radica, M., Welbanks, L., Espinoza, N., et al. 2023, MNRAS, 524, 835
- Ranjan et al. (2014) Ranjan, S., Charbonneau, D., Désert, J.-M., et al. 2014, ApJ, 785, 148
- Rathcke et al. (2021) Rathcke, A. D., MacDonald, R. J., Barstow, J. K., et al. 2021, AJ, 162, 138
- Rogers & Owen (2021) Rogers, J. G. & Owen, J. E. 2021, MNRAS, 503, 1526
- Rustamkulov et al. (2023) Rustamkulov, Z., Sing, D. K., Mukherjee, S., et al. 2023, Nature, 614, 659
- Sandford & Kipping (2017) Sandford, E. & Kipping, D. 2017, The Astronomical Journal, 154, 228
- Shulyak et al. (2020) Shulyak, D., Lara, L. M., Rengel, M., & Nèmec, N. E. 2020, A&A, 639, A48
- Sikora et al. (2023) Sikora, J., Rowe, J., Barat, S., et al. 2023, AJ, 165, 250
- Sing et al. (2016) Sing, D. K., Fortney, J. J., Nikolov, N., et al. 2016, Nature, 529, 59
- Sing et al. (2024) Sing, D. K., Rustamkulov, Z., Thorngren, D. P., et al. 2024, Nature, 630, 831
- Speagle (2020) Speagle, J. S. 2020, MNRAS, 493, 3132
- Suárez Mascareño et al. (2021) Suárez Mascareño, A., Damasso, M., Lodieu, N., et al. 2021, Nature Astronomy [arXiv:2111.09193]
- Swastik et al. (2021) Swastik, C., Banyal, R. K., Narang, M., et al. 2021, AJ, 161, 114
- Szabó & Kiss (2011) Szabó, G. M. & Kiss, L. L. 2011, ApJ, 727, L44
- Terquem & Papaloizou (2018) Terquem, C. & Papaloizou, J. C. B. 2018, Monthly Notices of the Royal Astronomical Society, 482, 530
- Thorngren & Fortney (2019) Thorngren, D. & Fortney, J. J. 2019, ApJ, 874, L31
- Tremblin et al. (2015) Tremblin, P., Amundsen, D. S., Mourier, P., et al. 2015, ApJ, 804, L17
- Tsai et al. (2023) Tsai, S.-M., Lee, E. K. H., Powell, D., et al. 2023, Nature, 617, 483
- Tsiaras et al. (2016) Tsiaras, A., Waldmann, I. P., Rocchetto, M., et al. 2016, The Astrophysical Journal, 832, 202
- Van Eylen et al. (2018) Van Eylen, V., Agentoft, C., Lundkvist, M. S., et al. 2018, MNRAS, 479, 4786
- Vazan et al. (2022) Vazan, A., Sari, R., & Kessel, R. 2022, The Astrophysical Journal, 926, 150
- Venot & Agúndez (2015) Venot, O. & Agúndez, M. 2015, Experimental Astronomy, 40, 469
- Wakeford & Dalba (2020) Wakeford, H. R. & Dalba, P. A. 2020, Philosophical Transactions of the Royal Society of London Series A, 378, 20200054
- Wakeford et al. (2013) Wakeford, H. R., Sing, D. K., Deming, D., et al. 2013, Monthly Notices of the Royal Astronomical Society, 435, 3481
- Wakeford et al. (2016) Wakeford, H. R., Sing, D. K., Evans, T., Deming, D., & Mandell, A. 2016, ApJ, 819, 10
- Welbanks et al. (2019) Welbanks, L., Madhusudhan, N., Allard, N. F., et al. 2019, ApJ, 887, L20
- Xue et al. (2024) Xue, Q., Bean, J. L., Zhang, M., et al. 2024, ApJ, 963, L5
- Yu et al. (2021) Yu, X., He, C., Zhang, X., et al. 2021, Nature Astronomy, 5, 822
- Zhou et al. (2020) Zhou, G., Quinn, S. N., Irwin, J., et al. 2020, The Astronomical Journal, 161, 2
- Zhou et al. (2017) Zhou, Y., Apai, D., Lew, B. W. P., & Schneider, G. 2017, The Astronomical Journal, 153, 243
@@ Appendix A Description of raw data reduction
We use the HST/WFC3 data reduction pipeline which has been presented in previous studies (Arcangeli et al. 2018, 2019, 2021; Jacobs et al. 2022; Barat et al. 2024). We do not change anything in the pipeline for the analysis of this paper. We start our analysis from the ‘ima.fits’ files downloaded from MAST archive. During each exposure, the detector is read multiple times, without flushing out the accumulated charge (non-destructive reads). We create sub-exposures by subtracting consecutive non-destructive reads for each exposure which are reduced separately. This technique has been widely applied in past observations with HST. We apply a wavelength dependent flat-field correction and flag bad pixels with data quality DQ=4, 32, or 512 by calwf3. We apply a local median filter to identify cosmic rays and clip pixels which are more than five median deviations from the median. To account for the dispersion direction drift of the spectrum (Wakeford et al. 2016) we use the first exposure of a visit as a template and shift the spectrum for an exposure along the dispersion direction to match the template. The maximum shift that we require is 0.1 pixels. We apply optimal extraction algorithm (Horne 1986) on each subexposure to maximize signal-to-noise ratio.
@@ Appendix B Light curve fitting
HST WFC3 light curves often exhibit strong systematics, such as ‘hook-like’ ramps and visit-long slopes (Berta et al. 2012; Deming et al. 2013; Ranjan et al. 2014; Tsiaras et al. 2016; Wakeford et al. 2016). Common practice involves discarding the first orbit due to significantly larger systematics (See Figure 1) compared to the rest of the visit (e.g. see, Wakeford et al. 2013; Arcangeli et al. 2018; Jacobs et al. 2022; Barat et al. 2024) from the light curve fitting. For the visit of planet b, the first orbit had been discarded (Barat et al. 2024). However, for the visit of planet c, we could not remove the first orbit of the visit from the analysis because the transit occurred two hours earlier than expected from the linear ephemeris (Feinstein et al. 2022) due to large TTVs and we could only secure one orbit prior to ingress.
@@
The first-orbit systematics are strongly wavelength dependent (Zhou et al. 2020, 2017). Therefore, the spectroscopic light curves cannot be de-trended using a common-mode approach, which had been previously applied to similar observations of V1298 tau b (Barat et al. 2024). Therefore, to handle these systematics, we used the physically motivated RECTE charge-trapping model (Zhou et al. 2017), which fits the first orbit ramp better than other known methods (e.g, Berta et al. 2012). We fixed the mid-transit time using simultaneous TESS observations (Feinstein et al. 2022). V1298 Tau shows strong rotational variability, with an average amplitude of 2 and a rotational period of 2.83 days (David et al. 2019b). We modeled it with a second-order polynomial for the out-of-transit baseline. See Appendix C for more discussion on the baseline modeling. The planetary transit signal was modeled using batman (Kreidberg 2015).
@@
For the white light curve (1.12-1.6 m), we fit the RECTE parameters and the baseline polynomial coefficients, along with the planet’s transit depth and semi-major axis. Other orbital parameters were fixed to literature values (Feinstein et al. 2022; Sikora et al. 2023; David et al. 2019a). The RECTE parameters constitute 4 parameters: the number of initially filled fast/slow charge traps and the number of fast/slow charge traps that are filled in between orbits. Due to the lack of ingress and egress data, our limb darkening precision was impaired. However, by using the limb darkening parameter retrieved from the observations of planet b (Barat et al. 2024), we were able to significantly reduce the uncertainties on the transit spectrum (Figure 9). V1298 Tau b and c have impact parameters (0.34 and 0.46 respectively) consistent within 1 of each other. Furthermore, Johnson et al. (2022) report a relatively low mutual inclination (019∘), indicating a planar orbital geometry of the V1298 Tau system. Therefore, both planets are likely to transit a similar part of the stellar photosphere with limb darkening during their primary transit. This highlights the advantage and uniqueness of studying multiple planets in the same system. For a discussion on the limb darkening of V1298 Tau see Section 4.2. We also independently measure the stellar density from the HST white light curves which is discussed in Section 4.1.
@@
The fit was performed using emcee (Foreman-Mackey et al. 2013). The posterior distribution from these fits is shown in Fig. 6 and the best-fit transit model and residuals, are shown in Figure 1 (upper panel).
@@
Spectroscopic light curves were analyzed similarly by dividing the stellar spectra into 7-pixel bins, resulting in 17 channels (identical to V1298 Tau b). We fixed the semi-major axis from the white light curve fits and fit the RECTE parameters and baseline for each bin, and the transit depth. We fix the wavelength dependent limb darkening coefficients to values derived in Barat et al. (2024). The de-trended spectroscopic light curves and residuals are shown in Fig. 7.
Appendix C Modeling the baseline